Determine if the statement is TRUE or FALSE. Justify your answer. All numbers under discussion are integers.


1.For each m ≥  1 and n ≥ 1, if mn is a multiple of 4, then m or n is a

multiple of 4.

2. For each m ≥ 1 and n ≥ 1, if mn is a multiple of 3, then m or n is a

multiple of 3.


1
Expert's answer
2021-09-24T13:16:28-0400

1. To show that this statement is false, we look for a counterexample

"m=2\geq1, n=2\geq1"


"mn=2(2)=4"

Since 4 is a multiple of 4, then "mn" is a multiple of 4. But "m=2" is not a multiple of 4, "n=2" is not a multiple of 4.

False.


2. "m" is not a multiple of 3 and "n" is not a multiple of 3.

Possible cases


"m=3a+1, n=3b+1,""mn=(3a+1)(3b+1)=9ab+3a+3b+1"

"mn" is not a multiple of 3. Contradiction.



"m=3a+1, n=3b+2,""mn=(3a+1)(3b+2)=9ab+6a+3b+2"

"mn" is not a multiple of 3. Contradiction.



"m=3a+2, n=3b+1,""mn=(3a+2)(3b+1)=9ab+3a+6b+2"

"mn" is not a multiple of 3. Contradiction.



"m=3a+2, n=3b+2,""mn=(3a+2)(3b+2)=9ab+6a+6b+4""=9ab+6a+6b+3+1"


"mn" is not a multiple of 3. Contradiction.

Therefore if mn is a multiple of 3, then "m\geq1" or "n\geq1" is a multiple of 3.

True.